OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (anonymous):

im totally lost

OpenStudy (anonymous):

@phi can u plz help, thanks

OpenStudy (anonymous):

do any of u know how to solve?

OpenStudy (ivancsc1996):

you need to take tHe referene triangle: THe Cot of 45 is te same as 1/tan(45) wiCH is tHe Hypotenuse/adjaent or \[\cot (45)=\frac{ 1 }{ \tan (45) }=\frac{ Hyp }{ adj }=\frac{ \sqrt{2} }{ 1 }=\sqrt{2}\]

OpenStudy (agent0smith):

@ivancsc1996 it should be cot = adj/opp, since tan is opp/adj.

OpenStudy (ivancsc1996):

OH you are rigHt

OpenStudy (anonymous):

cot = x/y = cos/sin? right?

OpenStudy (ivancsc1996):

tHen tHe answer is 1

OpenStudy (agent0smith):

Correct, cot = cos/sin

OpenStudy (anonymous):

and cos and sin are both 45 right?

OpenStudy (ivancsc1996):

no, tHey are BotH 1

OpenStudy (agent0smith):

No, cos45 = sin45 = sqrt2/2

OpenStudy (agent0smith):

When you divide them, you get 1.

OpenStudy (ivancsc1996):

YEaH wHat He said

OpenStudy (anonymous):

so its (sqrt2/2)/(sqrt2/2)

OpenStudy (anonymous):

which becomes sqrt/2 * 2/sqrt

OpenStudy (anonymous):

which then cancels to 1 right?

OpenStudy (ivancsc1996):

NO, it is \[\frac{ \frac{ 1 }{ \sqrt{2} } }{ \frac{ 1 }{ \sqrt{2} } }=\frac{ \cos 45 }{ \sin 45 }=\cot 45\]

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

@agent0smith could u explain

OpenStudy (agent0smith):

Yes, cot45 = 1, just like tan45=1.

OpenStudy (anonymous):

how do u get there though

OpenStudy (anonymous):

i need to see hoe u got there

OpenStudy (agent0smith):

Get cos and sine from the triangle (or use tan): cos45 = 1/sqrt2 sin45 = 1/sqrt2 cot45 = 1/tan45 tan45 = (1/sqrt2)/(1/sqrt2) = 1 so cot45 = 1/tan45 = 1/1 =

OpenStudy (anonymous):

why 1/sqrt 2

OpenStudy (agent0smith):

Because cos is adj/hyp... look at the triangle.

OpenStudy (anonymous):

why r the side lengths 1

OpenStudy (anonymous):

cos45 = 1 / sqrt2 sin45 = 1 / sqrt2 cot45 = cos / sin cot54 = (1/sqrt2)/(1/sqrt2) = 1/1 = 1

OpenStudy (anonymous):

is that a good explanation

OpenStudy (agent0smith):

The side lengths have to both be 1 with a right angle so that it's isosceles, and will have 45 degree angles.